# Investigating cannibalistic millisecond pulsar binaries using MESA : new constraints from pulsar spin and mass evolution

### v1

@@ 4.3 Pulsar spin evolution and sub-millisecond pulsars
Figure 11: Accreted matter versus final NS spins ($P^{f}_{\rm spin}$) in our simulated grid with $\beta=0.3$. We show the different initial masses as different colors (shown in the legend) and distinguish between the NSs that had $M^{f}_{NS}\geq 2.5\,{\rm M}_{\odot}$ (crosses) and those where $M^{f}_{NS}<2.5\,{\rm M}_{\odot}$ (circles). We also compare our results to the equilibrium spin period calculations from Tauris et al. (2012), done for $M^{i}_{NS}=1.3\,{\rm M}_{\odot}$ (solid blue curve) and $2.0\,{\rm M}_{\odot}$ (solid red curve). The gray dashed line shows the observed spin for PSR J0952–0607 (spin period of 1.41 ms).
@@
According to Figure 11, NSs in LMXBs and IMXBs accrete between 0.2 and $1.0\,{\rm M}_{\odot}$ to spin-up to millisecond periods ($1.38\rm\,ms\leq P^{f}_{\rm spin}\leq 30.0\rm\,ms$). NSs with initially higher masses ($\gtrsim 1.8\,{\rm M}_{\odot}$) accrete more than lighter NSs and hence, spin-up to a higher degree. Naturally, some of these heavy accretors could exceed the NS mass limit and enter the sub-MSP regime. Haskell et al. (2018) found that masses greater than $2.0\,{\rm M}_{\odot}$ have an upper limiting frequency of around 810 Hz (lower $P_{\rm spin}$ limit of 1.23 ms) depending on the NS equation of state, while lighter NSs could spin faster (up to 1200 Hz or 0.83 ms), inconsistent with observations. The spin at which an accreting NS would collapse into a BH also depends on the initial mass: for $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, it is at 1.38 ms and for $M^{i}_{\rm NS}=2.0\,{\rm M}_{\odot}$, it is at 2.66 ms. For a NS lighter than $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, the switch from NS to BH could occur at a spin lower than 1.38 ms. While our resulting spins are not as small as those by Haskell et al. (2018), we see a similar trend. For masses $\gtrsim 2.0\,{\rm M}_{\odot}$ the equation-of-state could change, leading to a collapse of the NS into a BH, or the formation of a hybrid star with sub-MSP spin (Gerlach, 1968; Glendenning & Kettner, 2000; Bejger et al., 2017). Liu et al. (2024) accounted for selection bias due to pulse broadening and suggested chances for Square Kilometer Array to observe sub-millisecond pulsar spins would be negligible. Since there are no current observations of any such sub-MSP star, sub-MSPs may not exist if those NSs have collapsed into BH.
4.4 Uncertainties in RB formation
@@
We find an evolutionary connection between UCXBs, RBs, and tiderrans (see Section 3). Our results for the formation of UCXBs are similar to other studies where they find UCXBs originating from LMXBs starting from close to $P_{\rm bif}$ (Podsiadlowski et al., 2002; van der Sluys et al., 2005b, a; Istrate et al., 2014). The idea that the Tidarren PSR J1311–3430 went through a UCXB phase was also hinted at by Chen et al. (2013) and. There is the fine-tuning problem in the formation of UCXBs. With the standard magnetic braking prescription (Skumanich, 1972; Verbunt & Zwaan, 1981; Rappaport et al., 1983) often used (also used in our study), there is a narrow range of orbital periods (that is near-$P_{\rm bif}$) and companion masses from which UCXBs can be produced (van der Sluys et al., 2005b, a; Istrate et al., 2014). Even though strong irradiation widens near-$P_{\rm bif}$ orbits and some of them might no longer be UCXBs, some UCXBs can still be reproduced (see $f_{\rm pulsar}=0.5$ in Figure 1).
@@
The dependence of $P_{\rm bif}$ on magnetic braking assumptions and the effects on LMXB evolution have been extensively studied (Podsiadlowski et al., 2002; Li et al., 2004; Ma & Li, 2009; Chen & Podsiadlowski, 2016; Deng et al., 2021; El-Badry et al., 2022; Echeveste et al., 2024), and it has been suggested that a stronger magnetic braking process is needed to form UCXBs (and consequently RBs) from a larger initial parameter space. In regards to spiders, Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021) focused on reproducing the observed BW properties and proposed a sustained magnetic braking prescription that will drive RLO; magnetic braking will not stop operating at $M_{\rm c}\sim 0.2\,{\rm M}_{\odot}$ and the companion has lost its radiative core. Gossage et al. (2023) compared single star and LMXB observations and found saturated magnetic braking to best reproduce the observations. Still the standard magnetic braking prescription can reproduce LMXBs with periods less than 1 hr to around 10 d, better than some other prescriptions (Echeveste et al., 2024; Gossage et al., 2023). Carrying out a population synthesis study of UCXBs would also tell us what fraction of LMXBs in a stellar population start RLO close to their $P_{\rm bif}$. Hence, to disentangle the various formation channels and types of magnetic braking prescriptions, a combined investigation of detailed simulations and population synthesis analysis would be required, which would further help us in understanding the formation of spiders.
@@ 5 Conclusions
Compact MSP binaries (or spider binaries) are good laboratories for studying binary evolution. We calculated the NS spin evolution and identified binaries that formed MSPs (using the criteria for NS spins: $P^{f}_{\rm pular}\leq 30$ ms). Assuming a model for an X-ray irradiated accretion disc and three different values of the accretion efficiency (0.0, 0.3, and 0.7), we ran simulation grids spanning four values of initial NS mass ($1.3\,{\rm M}_{\odot}$, $1.5\,{\rm M}_{\odot}$, $1.8\,{\rm M}_{\odot}$, and $2.0\,{\rm M}_{\odot}$), a range of initial companion masses ($1.0\,{\rm M}_{\odot}$ to $4.5\,{\rm M}_{\odot}$), and initial orbital periods (0.6 d to 4.0 d). Following the RLO phase, as the pulsar spins down, it loses rotational energy in the form of spin-down luminosity, ablating the outer envelope of the companion star and causing changes in the orbital evolution. We used four different values for the efficiency of this process (0.0, 0.05, 0.1, and 0.5) and presented the allowed initial parameter space to produce BWs, RBs, and huntsmen spiders. Comparing our simulated grids to the observed spider populations, we arrive at the following conclusions:
@@
  • With the assumption of unstable accretion disc due to X-ray irradiation during the mass-transfer phase, we need an accretion efficiency of 70% to reproduce the observed spins and pulsar masses of BW systems. High accretion efficiency ($>70\%$) also allows for the formation of the huntsmen spiders from a binary with a canonical NS.
@@
  • For BW formation, a high pulsar wind efficiency ($f_{\rm pulsar}=0.5$) are required to reproduce the full observed parameter space, although lower efficiencies ($0.50.0<f_{\rm pulsar}<$ 0.5) can still reproduce part of the systems. However, the tidarrens (J1311–3430, J1653–0159, and J0636+5128; BWs with orbital periods $\lesssim 1.0$ d and $0.01$M companions) require no irradiation. We also predict that J0636+5128 (similar to its neighbours) should not have any hydrogen on its companion’s surface as the companion loses its hydrogen envelope through stable mass transfer. The binary tracks passing through the tidarren region also pass through the RB region, connecting BWs and RBs as part of the same initial population. These binaries start RLO at near-$P_{\rm bif}$ and pass through an UCXB phase. Their evolution is dependent on the magnetic braking prescription used.
@@
  • With high pulsar wind irradiation ($f_{\rm pulsar}=0.5$), the orbits of the near-$P_{\rm bif}$ binaries expand drastically, covering most of the RB region. This is true for the companion masses $\lesssim 1.5\,{\rm M}_{\odot}$, the near-$P_{\rm bif}$ binaries are the main formation channel for RBs. Hence, RBs are subject to the fine-tuning problem seen in UCXBs, where only a narrow range of initial parameter space can reproduce the observations.
@@
  • We identified the initial parameter space for diverging NS binaries to spin-up as MSPs. These systems would form the small population of huntsmen spiders that have been observed. With fully conservative accretion (assuming an X-ray irradiated accretion disc) a NS in diverging orbits with initial mass $1.3\,{\rm M}_{\odot}$ gains a maximum of $0.55\,{\rm M}_{\odot}$. If a huntsman pulsar is discovered to have a pulsar mass $\gtrsim 2.0\,{\rm M}_{\odot}$, most likely it was formed more massive than $1.3\,{\rm M}_{\odot}$.
@@
  • We identified the prior evolution of the pulsars: BW PSR J0952–0607 and huntsman 1FGL J1417.7–440, by matching their observed spins and orbital properties to binary tracks. For PSR J0952–0607, high irradiation ($f_{\rm pulsar}=0.5$) after case A RLO is required to reproduce its estimated orbital properties. For 1FGL J1417.7–440, the decreased expansion of the orbit due to pulsar wind irradiation ($f_{\rm pulsar}=0.1$) matches the observations well, though irradiation is not required to reproduce the orbital properties of these diverging NS binaries.
@@
  • NSs in LMXBs/IMXBs accrete between 0.2 and $1.0\,{\rm M}_{\odot}$ to spin-up to MSP spins. For NSs to attain sub-MSP spins, the mass accreted is large enough to collapse it into a BH ($\gtrsim 1.0\,{\rm M}_{\odot}$ for $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$). Additionally, the spin at which an accreting NS would enter the sub-MSP regime (and collapse) increases with increasing mass, going from 1.38 ms to 2.66 ms when $M^{i}_{\rm NS}$ goes from $1.3\,{\rm M}_{\odot}$ to $2.0\,{\rm M}_{\odot}$. This could explain why sub-MSPs are not observed and why the fastest spinning BW spins at 1.4 ms. With $\beta=0.3$, we can reproduce the fastest spinning BW currently observed PSR J0952–0607 (with observed spin 1.41 ms and pulsar mass of $2.35\,{\rm M}_{\odot}$). Hence, fast spinning NSs are key in identifying the boundary between NSs and BHs.
@@
  • Around 50 to 60% of our simulated spiders ended up becoming dynamically unstable due to mass transfer initiated from a fully convective companion. These binaries would go through a common envelope phase, the surviving binaries for which would be isolated spinning NSs. These could explain the observations of isolated MSPs.
@@
Since this study involved running detailed MESA  simulations with a fixed set of initial parameters, we do not know the formation probability of these types of systems and how it would compare to the observed populations. In order to calculate the formation rates of spiders we would need to carry out a population synthesis study. Using large-scale parameter studies, we can better constrain certain physical parameters governing binary evolution. These include processes affecting the evolution prior to the XRB phase (like the supernova mechanism) and more directly involved physical processes (like magnetic braking and pulsar wind irradiation). Since LMXBs have very unstable accretion discs causing outbursts that would affect pulsar spin-up (for instance, due to X-ray irradiation during RLO; Dubus et al., 1999), unsteady or transient accretion must be further investigated in detail. While we do include a model to account for these effects that would cause non-conservative mass transfer, these instabilities and their effects on the observed populations are highly uncertain and require further investigation. Finally, this study was conducted at solar metallicity, because of the substantial population of Galactic spiders that have been observed. However, other metallicities should also be investigated to account for the spiders that have been observed in globular clusters with metallicities below solar (e.g., PSR J1953+1844 in the globular cluster M71 at 1/10th solar metallicity; Pan et al., 2023).
@@ Acknowledgements.
The authors thank Thomas Tauris and Hai-Liang Chen for their useful discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580.
@@ Appendix A Numerical stability of MESA tracks
Figure 12: Stellar luminosity ($L$) versus effective temperature ($T_{\rm eff}$) of a binary from the simulated grids, with an initial companion mass of 4.5 M, initial NS mass of 2.0 M, and initial orbital period of 1.4 d, for different maximum opacity values ($\kappa_{\rm max}$). The solid blue curve is the failed binary. The solid green, dashed red, and dashed purple curves are with $\kappa_{\rm max}$ value of 10.0, 5.0, and 0.5, respectively. We compare the tracks to a neighboring track (that ended without any numerical instability) in solid orange.
@@
Fragos et al. (2022, Section 5.2) list a wide range of criteria for successful termination of the MESA  track, (1) the age of the companion to the NS exceeds the Hubble age (13.8 Gyr), (2) the companion to the NS becomes a WD, and (3) the binary enters a common envelope. The binaries that did not satisfy the criteria for successful termination (mentioned in Section 2.1), were re-run with a limitation on the maximum radiative opacity ($\kappa_{\rm max}$) of the companion. The initial failure rate was 16.5% in the entire initial simulated grid. Since the stellar opacity influences the internal stellar structure, it prevents the star properties from fluctuating too drastically for MESA  to handle. The re-run $\kappa_{\rm max}$ value used by Fragos et al. (2022) is 0.5 $\rm cm{{}^{2}}\,g^{-1}$, which brought down our failure rate to 0.27%. However, we find this value to be too restrictive for some stars in our initial parameter space, resulting in a significant deviation in the resulting tracks from the successful tracks in close vicinity. This can be seen in Figure 12, where we show a binary (with $\beta=0.0$) having an initial companion mass of 4.5 M, initial NS mass of 2.0 M, and initial orbital period of 1.2 d, that failed due to numerical instability. We also show a neighbouring binary that avoided numerical instabilities with 4.5 M, initial NS mass of 2.0 M, and initial orbital period of 1.4 d. Decreasing the value of $\kappa_{\rm max}$ to 0.5 results in the stellar properties deviating significantly from previous reruns, from the very beginning of the evolution. The binary, that previously should have undergone a dynamical instability, with a strict $\kappa_{max}$ does not have a high mass-transfer rate and ends as a WD. Hence, to keep the tracks as similar to the original run as possible (while avoiding numerical instabilities), as well as similar in behaviour to their close neighbours, we re-run the failed binaries in two batches, first we used a $\kappa_{\rm max}$ of 10 $\rm cm{{}^{2}}\,g^{-1}$. Then we run the binaries that failed again, with a $\kappa_{\rm max}$ of 5 $\rm cm{{}^{2}}\,g^{-1}$. Our final failure rate was 9.19%, which is a lot higher than the case of a strict $\kappa_{\rm max}$, however we compromised by checking the final properties of the binaries and if they seemed to have an MSP-like spin for the NS ($\leq 30$ ms), we considered the binary successful.

### v2

@@ 4.3 Pulsar spin evolution and sub-millisecond pulsars
Figure 11: Accreted mass versus final NS spins ($P^{f}_{\rm spin}$) in our simulated grid with $\beta=0.3$. We show the different initial masses as different colors (shown in the legend) and distinguish between the NSs that had $M^{f}_{NS}\geq 2.5\,{\rm M}_{\odot}$ (crosses) and those where $M^{f}_{NS}<2.5\,{\rm M}_{\odot}$ (circles). We also compare our results to the equilibrium spin period calculations from Tauris et al. (2012), done for $M^{i}_{NS}=1.3\,{\rm M}_{\odot}$ (light blue curve) and $2.0\,{\rm M}_{\odot}$ (dark orange curve). The black dotted line shows the observed spin for PSR J0952–0607 (spin period of 1.41 ms) and the gray dotted line shows the upper limit for MSPs considered (30 ms).
@@
According to Figure 11, NSs in LMXBs and IMXBs accrete between 0.2 and $1.0\,{\rm M}_{\odot}$ to spin up to millisecond periods ($1.38\rm\,ms\leq P^{f}_{\rm spin}\leq 30.0\rm\,ms$). The NSs with initially higher masses ($\gtrsim 1.8\,{\rm M}_{\odot}$) accrete more than lighter NSs and hence spin up to a higher degree. Naturally, some of these heavy accretors could exceed the NS mass limit and enter the sub-MSP regime. Accretion induced collapse (AIC) of massive NSs was also studied by Chen et al. (2023) where they found inefficient RLO required initial NS masses of at least 2.16 ${\rm M}_{\odot}$in order for AIC to occure. Haskell et al. (2018) found that masses greater than $2.0\,{\rm M}_{\odot}$ have an upper limiting frequency of around 810 Hz (lower $P_{\rm spin}$ limit of 1.23 ms) depending on the NS equation of state, while lighter NSs could spin faster (up to 1200 Hz or 0.83 ms), inconsistent with observations. The spin at which an accreting NS would collapse into a BH also depends on the initial mass: for $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, it is at 1.38 ms and for $M^{i}_{\rm NS}=2.0\,{\rm M}_{\odot}$, it is at 2.66 ms. For an NS lighter than $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, the switch from NS to BH could occur at a spin lower than 1.38 ms. While our resulting spins are not as small as those by Haskell et al. (2018), we see a similar trend. For masses $\gtrsim 2.0\,{\rm M}_{\odot}$ the equation-of-state could change, leading to a collapse of the NS into a BH, or the formation of a hybrid star with sub-MSP spin (Gerlach, 1968; Glendenning & Kettner, 2000; Bejger et al., 2017). Liu et al. (2024) accounted for selection bias due to pulse broadening and suggested chances for Square Kilometer Array to observe sub-millisecond pulsar spins would be negligible. Since there are no current observations of any such sub-MSP star, sub-MSPs may not exist if those NSs have collapsed into BH.
4.4 Uncertainties in redback formation
@@
We find an evolutionary connection between UCXBs, RBs, and tiderrans (see Section 3). Our results for the formation of UCXBs are similar to other studies where they find UCXBs originating from LMXBs starting from close to $P_{\rm bif}$ (Podsiadlowski et al., 2002; van der Sluys et al., 2005b, a; Istrate et al., 2014; He et al., 2019). The idea that the Tidarren PSR J1311–3430 went through a UCXB phase was also hinted at by Chen et al. (2013) and. There is the fine-tuning problem in the formation of UCXBs. With the standard magnetic braking prescription (Skumanich, 1972; Verbunt & Zwaan, 1981; Rappaport et al., 1983) often used (also used in our study), there is a narrow range of orbital periods (that is near-$P_{\rm bif}$) and companion masses from which UCXBs can be produced (van der Sluys et al., 2005b, a; Istrate et al., 2014). Even though strong irradiation widens near-$P_{\rm bif}$ orbits and some of them might no longer be UCXBs, some UCXBs can still be reproduced (see $f_{\rm pulsar}=0.5$ in Figure 1).
@@
The dependence of $P_{\rm bif}$ on magnetic braking assumptions and the effects on LMXB evolution have been extensively studied (Podsiadlowski et al., 2002; Li et al., 2004; Ma & Li, 2009; Chen & Podsiadlowski, 2016; Deng et al., 2021; El-Badry et al., 2022; Echeveste et al., 2024), and it has been suggested that a stronger magnetic braking process is needed to form UCXBs (and consequently RBs) from a larger initial parameter space. Chen et al. (2021) tested the magnetic braking prescription suggested by Van & Ivanova (2019) in the formation of binary MSPs and UCXBs. They found that while efficient magnetic braking does increase the allowed initial parameter space to form UCXBs, wide-orbit binary MSPs are not reproduced. In regards to spiders, Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021) focused on reproducing the observed BW properties and proposed a sustained magnetic braking prescription that will drive RLO; magnetic braking will not stop operating at $M_{\rm c}\sim 0.2\,{\rm M}_{\odot}$ and the companion has lost its radiative core. Gossage et al. (2023) compared MESA simulations with single star and LMXB observations and also found saturated magnetic braking to best reproduce both types of sources. Still the standard magnetic braking prescription can reproduce LMXBs with periods less than one hour to around ten days, better than some other proposed prescriptions (Echeveste et al., 2024; Gossage et al., 2023). Carrying out a population synthesis study of UCXBs would also tell us what fraction of LMXBs in a stellar population start RLO close to their $P_{\rm bif}$. Hence, to disentangle the various formation channels and types of magnetic braking prescriptions, a combined investigation of detailed simulations and population synthesis analysis would be required, which would further help in understanding the formation of spiders.
@@ 5 Conclusions
Compact MSP binaries (or spider binaries) are good laboratories for studying binary evolution. In this work, we calculated the NS spin evolution and identified binaries that formed MSPs (using the criteria for NS spins: $P^{f}_{\rm pular}\leq 30$ ms). Assuming a model for an X-ray irradiated accretion disk and three different values of the accretion efficiency (0.0, 0.3, and 0.7), we ran simulation grids spanning four values of initial NS mass ($1.3\,{\rm M}_{\odot}$, $1.5\,{\rm M}_{\odot}$, $1.8\,{\rm M}_{\odot}$, and $2.0\,{\rm M}_{\odot}$), a range of initial companion masses ($1.0\,{\rm M}_{\odot}$ to $4.5\,{\rm M}_{\odot}$), and initial orbital periods (0.6 d to 4.0 d). Following the RLO phase, as the pulsar spins down, it loses rotational energy in the form of spin-down luminosity, ablating the outer envelope of the companion star and causing changes in the orbital evolution. We used four different values for the efficiency of this process (0.0, 0.05, 0.1, and 0.5) and presented the allowed initial parameter space to produce BWs, RBs, and huntsmen spiders. After comparing our simulated grids to the observed spider populations, we arrived at the following conclusions:
@@
  • With the assumption of an unstable accretion disk due to X-ray irradiation during the mass-transfer phase, an accretion efficiency of at least 70% is needed in order to reproduce the observed spins and pulsar masses of BW systems. High accretion efficiency ($>70\%$) also allows for the formation of the huntsmen spiders from a binary with a canonical NS in a widening orbit during RLO.
@@
  • For BW formation, a high pulsar wind efficiency ($f_{\rm pulsar}=0.5$) is required to reproduce the full observed parameter space, although lower efficiencies ($0.50.0<f_{\rm pulsar}<$ 0.5) can still reproduce part of the systems. However, the tidarrens (J1311–3430, J1653–0159, and J0636+5128; BWs with orbital periods $\lesssim 1.0$ d and $0.01$M companions) require no irradiation. We also predict that J0636+5128 (similar to its neighbors) should not have any hydrogen on its companion’s surface, as the companion loses its hydrogen envelope through stable mass transfer. The binary tracks passing through the tidarren region also pass through the RB region, connecting BWs and RBs as part of the same initial population. These binaries start RLO at near-$P_{\rm bif}$ and pass through an UCXB phase. Their evolution is dependent on the magnetic braking prescription used.
@@
  • With high pulsar wind irradiation ($f_{\rm pulsar}=0.5$), the orbits of the near-$P_{\rm bif}$ binaries expand drastically, covering most of the RB region. This is true for the companion masses $\lesssim 1.5\,{\rm M}_{\odot}$. The near-$P_{\rm bif}$ binaries are the main formation channel for RBs. Hence, RBs are subject to the fine-tuning problem seen in UCXBs, where only a narrow range of values for the initial orbital parameter space, with standard magnetic braking, can reproduce the observations.
@@
  • We identified the initial parameter space for diverging NS binaries to spin up as MSPs. These systems would form the small population of huntsmen spiders that have been observed. With fully conservative accretion (assuming an X-ray irradiated accretion disk), an NS in diverging orbits with an initial mass of $1.3\,{\rm M}_{\odot}$ gains a maximum of $0.55\,{\rm M}_{\odot}$. If a huntsman pulsar is discovered to have a pulsar mass $\gtrsim 2.0\,{\rm M}_{\odot}$, it was most likely more massive than $1.3\,{\rm M}_{\odot}$ when formed.
@@
  • We identified the prior evolution of the pulsars BW PSR J0952–0607 and huntsman 1FGL J1417.7–440 by matching their observed spins and orbital properties to binary tracks. For PSR J0952–0607, high irradiation ($f_{\rm pulsar}=0.5$) after case A RLO is required to reproduce its estimated orbital properties. For 1FGL J1417.7–440, the decreased expansion of the orbit due to pulsar wind irradiation ($f_{\rm pulsar}=0.1$) matches the observations well, though irradiation is not required to reproduce the orbital properties of these diverging NS binaries.
@@
  • The NSs in LMXBs and IMXBs accrete between 0.2 and $1.0\,{\rm M}_{\odot}$ to spin up to MSP spins. For NSs to attain sub-MSP spins, the mass accreted is large enough to collapse it into a BH ($\gtrsim 1.0\,{\rm M}_{\odot}$ for $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$). Additionally, the spin at which an accreting NS would enter the sub-MSP regime (and collapse) increases with increasing mass, going from 1.38 ms to 2.66 ms when $M^{i}_{\rm NS}$ goes from $1.3\,{\rm M}_{\odot}$ to $2.0\,{\rm M}_{\odot}$. This could explain why sub-MSPs are not observed and why the fastest spinning BW spins at 1.4 ms. With $\beta=0.3$, we can reproduce the fastest spinning BW currently observed PSR J0952–0607 (observed spin of 1.41 ms and pulsar mass of $2.35\,{\rm M}_{\odot}$). Hence, fast spinning NSs are key in identifying the boundary between NSs and BHs.
@@
  • Around 50 to 60% of our simulated spiders ended up becoming dynamically unstable due to mass transfer initiated from a fully convective companion. These binaries would go through a common envelope phase, and the surviving binaries would be isolated spinning NSs. These binaries could explain the observations of isolated MSPs.
@@
Since this study involved running detailed MESA  simulations with a fixed set of initial parameters, we did not know the formation probability of these types of systems and how it would compare to the observed populations. In order to calculate the formation rates of spiders, we would need to carry out a population synthesis study. Using large-scale parameter studies, we can better constrain certain physical parameters governing binary evolution. These include processes affecting the evolution prior to the XRB phase (e.g., the supernova mechanism) and more directly involved physical processes (e.g., magnetic braking and pulsar wind irradiation). Since LMXBs have very unstable accretion disks that cause outbursts that would affect pulsar spin-up (for instance, due to X-ray irradiation during RLO; Dubus et al., 1999), unsteady or transient accretion must be further investigated in detail. While we do include a model to account for these effects that would cause non-conservative mass transfer, these instabilities and their effects on the observed populations are highly uncertain and require further investigation. Finally, this study was conducted at solar metallicity because of the substantial population of Galactic spiders that have been observed. However, other metallicities should also be investigated to account for the spiders that have been observed in globular clusters with metallicities below solar (e.g., PSR J1953+1844 in the globular cluster M71 at 1/10th solar metallicity; Pan et al., 2023).
@@ Acknowledgements.
The authors thank the anonymous referee for their constructive comments that helped improve the manuscript. The authors thank Thomas Tauris, Hai-Liang Chen, and Raphaël Mignon-Risse for their useful discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580.
@@ Appendix A Numerical stability of MESA tracks
Figure 12: Stellar luminosity ($L$) versus effective temperature ($T_{\rm eff}$) of a binary from the simulated grids with an initial companion mass of 4.5 ${\rm M}_{\odot}$, initial NS mass of 2.0 ${\rm M}_{\odot}$, and initial orbital period of 1.4 d for different maximum opacity values ($\kappa_{\rm max}$). The solid blue curve is the failed binary. The solid green, dashed red, and dashed purple curves are with $\kappa_{\rm max}$ value of 10.0, 5.0, and 0.5, respectively. We compare the tracks to a neighboring track (without any numerical instability) in solid orange.
@@
Fragos et al. (2022, Section 5.2) list a wide range of criteria for successful termination of the MESA  track, (1) the age of the companion to the NS exceeds the Hubble age (13.8 Gyr), (2) the companion to the NS becomes a WD, and (3) the binary enters a common envelope. The binaries that did not satisfy the criteria for successful termination (mentioned in Section 2.1), were rerun with a limitation on the maximum radiative opacity ($\kappa_{\rm max}$) of the companion. The initial failure rate was 16.5% in the entire initial simulated grid. Since the stellar opacity influences the internal stellar structure, it prevents the star properties from fluctuating too drastically for MESA  to handle. The rerun $\kappa_{\rm max}$ value used by Fragos et al. (2022) is 0.5 $\rm cm{{}^{2}}\,g^{-1}$, which brought down our failure rate to 0.27%. However, we find this value to be too restrictive for some stars in our initial parameter space, resulting in a significant deviation in the resulting tracks from the successful tracks in close vicinity. This can be seen in Figure 12, where we show a binary (with $\beta=0.0$) having an initial companion mass of 4.5 ${\rm M}_{\odot}$, an initial NS mass of 2.0 ${\rm M}_{\odot}$, and an initial orbital period of 1.2 d, which failed due to numerical instability. We also show a neighboring binary that avoided numerical instabilities with 4.5 ${\rm M}_{\odot}$, initial NS mass of 2.0 ${\rm M}_{\odot}$, and initial orbital period of 1.4 d. Decreasing the value of $\kappa_{\rm max}$ to 0.5 results in the stellar properties deviating significantly from previous reruns, from the very beginning of the evolution. The binary that previously should have undergone a dynamical instability with a strict $\kappa_{max}$ does not have a high mass-transfer rate and ends as a WD. Hence, to keep the tracks as similar to the original run as possible (while avoiding numerical instabilities) as well as similar in behavior to their close neighbors, we reran the failed binaries in two batches. First we used a $\kappa_{\rm max}$ of 10 $\rm cm{{}^{2}}\,g^{-1}$, and then we ran the binaries that failed again with a $\kappa_{\rm max}$ of 5 $\rm cm{{}^{2}}\,g^{-1}$. Our final failure rate was 9.19%, which is a lot higher than the case of a strict $\kappa_{\rm max}$; however, we compromised by checking the final properties of the binaries, and if they seemed to have an MSP-like spin for the NS ($\leq 30$ ms), we considered the binary successful.